In an arithmetic sequence the seventh term is 9 times the first and the fourth
term is less than the fifth term by 4: Find the first term and the common
difference.

Answer by mananth(16946)

In an arithmetic sequence the seventh term is 9 times the first and the fourth
term is less than the fifth term by 4: Find the first term and the common
difference.
tn = a+(n-1)d
t7=a+6d
a+6d = 9a
8a=6d
4a=3d
---
T4 = a+3d
t5=a+4d


a+4d-(a+3d)=4
a+4d-a-3d=4
d=4
4a=3d
d=4
4a=12
a=3
a=3, d=4